OpenStudy (anonymous):

Find the lowest common denominator that would eliminate the fractions in this equation. y over 3+y over 4=5 over 6 (y/3+y/4=5/6)

OpenStudy (anonymous):

This shouldn't be too hard. Common denominator, convert to proportion, and solve.\[y/3+y/4=5/6\implies \frac{7y}{12}=\frac{5}{6}\implies 42y=60 \implies y=\frac{10}{7}\]

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

What don't you understand?

OpenStudy (anonymous):

Im only 10, i just wanted to know the answer because my friend was doing it earlier

OpenStudy (anonymous):

Good night.

OpenStudy (campbell_st):

the common denominator is 12... so multiply every term by 12 \[12 \times y/3 + 12 \times y/4 = 12 \times 5\] this can be simplified to look at the 1st term \[12 \times y/3 = (3 \times4\times y)3 = 4y\] the 3s cancel out

OpenStudy (campbell_st):

the next term is \[12y/4 = (4 \times 3\times y)/4\] the 4s cancel and 12 x 5 = 60 so the problem becomes 4y + 3y = 60 or 7y = 60

OpenStudy (campbell_st):

the common denominator of 12 comes from 3 x 4 the bottom of each fraction

OpenStudy (lgbasallote):

i taught this to you before right? list your denominators 3, 4, 6 break them down 3 -> 3 4 -> 2(2) 6 ->(2)(3) your LCD needs to contain ALL of these..so your LCD will be (3)(2)(2) = 12 check does it contain 3? \(\checkmark\) does it contain 4 or (2)(2)? \(\checkmark\) does it contain 6 or (2)(3)? \(\checkmark\) so 12 is the LCD

OpenStudy (anonymous):

When you multiply the RHS by 12, you get ten, not sixty.

OpenStudy (anonymous):

Otherwise, your explanation is great.

OpenStudy (anonymous):

Thanks @igbasalotte

OpenStudy (anonymous):

@igbasalotte will you fan me so i can send you messages

OpenStudy (lgbasallote):

it's lgbasallote